# Data-Driven Chance-Constrained Multiple-Choice Knapsack Problem: Model, Algorithms, and Applications

### v1

@@ C ET ECL RCL FSR C ET ECL RCL FSR
TABLE VI: Performance on benchmarks of APP. Due to the introduction of retransmission mechanisms, there is a significant difference in the results between LAB and APP.
Benchmark small $T_{\max}$ large $T_{\max}$
C ET ECL RCL FSR C ET ECL RCL FSR
ss2-DDALS(O) 5.2$\pm$0.0 307$\pm$7.6 99.2$\pm$0.0 97.3$\pm$0 0 5.2$\pm$0.0 291$\pm$10.6 99.6$\pm$0.0 98.1$\pm$0.0 0
ss2-DDALS(V1) 5.5$\pm$0.1 307$\pm$7.6 99.8$\pm$0.0 97.5$\pm$0.3 0 5.2$\pm$0 291$\pm$10.1 99.6$\pm$0.0 98.1$\pm$0.0 0
ss4-EDA 26.1$\pm$0.0 2005.7$\pm$2.3 100$\pm$0 90.1$\pm$0 0 16.4$\pm$0.0 1602.7$\pm$2.1 99.9$\pm$0.0 80.3$\pm$0.0 0
ss4-DDALS(O) 26.1$\pm$0.0 1988.2$\pm$30.4 1$\pm$0 90.1$\pm$0 0 16.7$\pm$0.9 1579.4$\pm$76.8 99.8$\pm$0.2 81.8$\pm$4.4 0
ss4-DDALS(V1) 26.1$\pm$0.0 1988.2$\pm$30.4 1$\pm$0 90.1$\pm$0 0 16.7$\pm$1.0 1579.4$\pm$ 76.8 99.9$\pm$0.0 81.5$\pm$3.6 0
ss4-DDALS(V2) 26.1$\pm$0.0 1988.2$\pm$30.4 1$\pm$0 90.1$\pm$0 0 18.0$\pm$1.9 1579.4$\pm$76.8 99.9$\pm$0.1 86.1$\pm$7.0 0
In the case of LAB scenarios, the results of ss3, ss4 and ls4 are shown in Table V. We can intuitively see that the DDALS algorithm outputs the lowest cost solution across all test benchmarks, while Greedy needs only less than 10 times of evaluation but results in the highest cost. The feasibility ratio of all algorithms is close to 1, indicating that our solution evaluation algorithm only needs a very few samples to obtain an evaluation that is close to the true confidence level for small-scale scenarios and 500 samples for large-scale scenarios. From another perspective, for other variations, the original version of DDALS can obtain solutions with lower cost than other variations while guaranteeing a feasible ratio of 1, indicating that the solution evaluation is accurate.
@@
There are significant changes in the scenario of APP, shown in Table V. In small-scale benchmark tests, only the Greedy achieves a feasibility ratio of 1, while the feasibility ratio of other algorithms is almost zero. Although DDALS can obtain lower cost solutions, all of them are found to be infeasible. Such results are practically useless. The main reason is due to the retransmission mechanism in APP, which introduces discontinuities in the item distribution that pose a significant challenge to the solution evaluation algorithm. In a random sampling case, small samples may easily miss information regarding exceptional jumps, leading to performance decline of solution evaluation. Greedy can achieve higher feasibility ratios, but it sacrifices the optimality of the solutions. For the results of large-scale problems, although the difficulty of search increased with the increase of scale, DDALS can still find feasible solutions achieving the lowest cost. Note that in the result of ls4 with large $T_{\max}$, the feasibility ratio of DDALS(O) is 0.3 while that of DDALS(V3) is 1. Given that the average quality of the solutions generated by DDALS(V3) decreases by only 3% relative to the original DDALS, the feasibility ratio increased by 233%. It shows the effectiveness and necessity of filtering the solutions in complex scenarios. Moreover, increasing the sample size $L$ can improve the evaluation performance, which can help capture more information and improve the accuracy of solution evaluation.
@@ VI-D Effectiveness of Components in DDALS
We conduct an ablation study to further assess the effectiveness of four components of DDALS, including CP, LSS, Degrade and FSS. To independently assess their effectiveness, four different variants of DDALS are proposed and tested on benchmarks of LAB:
@@
    r-CP DDALS: It uses random strategy to select the item in each class instead of utilizing CP. Note that if the initial randomly generated is infeasible solution, query the current solution for each class of item and re-generate randomly in each class of items less than the weight of the current item.
@@
    no-LSS DDALS: The LSS procedure in the iteration process is removed from DDALS.
@@
    no-Degrade DDALS: The degrade procedure in the iteration process is removed.
@@
    no-FSS DDALS: The FSS procedure is removed.
For all the above variants, we compute the performance degradation ratio(PDR) by computing $(C_{2}-C_{1})/C1$ for each benchmark, where $C_{1}$ and $C_{2}$ are the average cost of solution obtained by the variant and original across 10 independent runs. The component being examined is useful for a specific LAB benchmark if the result shows that PDR $>0$. Table VII presents the average PDR averaged on each benchmark of LAB. Based on the results, it can be concluded that none of the examined components has a negative effect on the performance of DDALS. In most cases, these components make LS perform better. Constructing utility for items is helpful for DDALS to obtain a good initial solution on average. LSS provides some kind of exploration, but a negative value on ls3 implies that due to the large size of the sample, the algorithm may still not converge after 30 iterations of running. The operation of degradation has the most significant contribution to the performance of LS, since it greatly improves the problem of the algorithm, as it helps the algorithm to escape local optima. A deeper search is necessary, while FSS has a slightly higher PDR than LSS. However, with the increase of the benchmark’s scale, the contribution of FSS seems to decrease, implying that more efficient and exploratory methods are needed to cope with larger-scale problems.
TABLE VII: Average PDR on each benchmark of LAB. The highest PDR is highlighted through bolding and being underlined.
Benchmark r-CP no-LSS no-Degrade no-FSS
@@ VI-E Acceleration Performance of Accelerated Monte Carlo
The acceleration effect of Accelerated Monte Carlo (AMC) evaluation is validated. We randomly selected 10,000 solutions for evaluation, using regular Monte Carlo (MC) and AMC respectively. After recording the running time required for each 1000 evaluations, the result is shown in Figure 6.
The results show that the time consumed increases essentially linearly as the number of evaluation times increases. For different $T_{\max}$ settings, MC takes essentially the same amount of time since the number of simulations needed does not vary with $T_{\max}$, while AMC can obtain better acceleration compared to MC when $T_{\max}$ is small. For example, compared with the running time $526.601s$ of MC, AMC takes $458.449s$ to finish the evaluation when $T_{\max}=11$, while the number of fast-filtered solution is 10000. When $T_{\max}=23$, it takes AMC about $471.395s$, although the number of pre-fast filtered is 9718. As $T_{\max}$ increases, the difficulty of the chance constraint decreases, resulting in the increase of the percentage of feasible solutions in the whole solution space. Thus, the number of solutions that can be fast pre-filtered by sequences decreases, which leads to a decrease in the acceleration effect. The accelerated mechanism may have a negative impact on the overall efficiency of the evaluation as $T_{\max}$ becomes large. When $T_{\max}=27$, AMC and MC take almost the same time, while when $T_{\max}=35$, AMC needs to take $807.435s$ to finish rather than about $500s$ of MC. In conclusion, when applying the accelerated mechanism, a pre-assessment of whether the constraints are tight enough is necessary to determine whether the constraints are tight enough.
Figure 6: Acceleration performance of accelerated Monte Carlo on LAB-ls1. The mechanism has a positive effect when the constraints are tight, but has a negative impact on the overall efficiency as $T_{\max}$ becomes large.
VI-F Performance of Other Works
We also attempted to adjust two similar algorithms, Ji’s [15] and Yang’s [34], to solve our problem (see Appendix 4 for modification details). Due to the incapability of dynamic programming and mixed integer programming for solving large-scale DDCCMCKP, we conducted experiments on small-scale benchmarks to evaluate their performance. Results are shown in Table VIII, and we observed that the majority of the results were either ”infeasible” or ”fail”. ”fail” means that the algorithms fail to output any solution, while ”infeasible” means that the optimal solution is infeasible. The value of ”Feasibility” column is the total cost of solution. For Yang’s method, the poor performance was mainly due to the significant difference between the actual distribution and the Gaussian distribution, and the insufficient data to characterize an unknown distribution using sample mean and variance, leading to incorrect iteration. In Ji’s work, a uniform distribution was used and tested. The performance of their method is poor since it could not handle the long-tailed distribution of the candidate solutions in our problem and led to suboptimal performance.
TABLE VIII: Result of Ji’s [15] and Yang’s [34]
Benchmark Ji Yang
Time(s) Feasiblility Time(s) Feasiblility
LAB-$ss1$ fail 1.0 14.5
LAB-$ss2$ 0.4 infeasible - fail
LAB-$ss3$ 0.8 infeasible 1.4 21.5
LAB-$ss4$ 16 infeasible 43.2 17.4
APP-$ss1$ fail 0.1 infeasible
APP-$ss2$ 0.6 infeasible - fail
APP-$ss3$ fail 0.2 infeasible
APP-$ss4$ fail 13.5 infeasible
@@ VII Conclusion
In this work, derived from practical applications, we study a new variant of MCKP called DDCCMCKP. To address this specific problem, a novel DDALS algorithm has been introduced. Based on local search framework, DDALS enables data-driven solution evaluation for the problem, and incorporates a surrogate weight to ensure the optimality of solutions. Moreover, a solution filtering mechanism is introduced to improve the feasibility of the output solutions. Two sets of benchmarks for DDCCMCKP have been proposed, one of which simulates the real business requirement of 5G E2E network scenario for a communication company. The experimental results demonstrate that DDALS is able to solve DDCCMCKP effectively and outperforms other baseline methods. Although the APP benchmark imposes higher demands on problem-solving abilitiy, the performance of DDALS is not satisfactory in small-scale problems. However, DDALS can solve DDCCMCKP well in most of the scenarios and maintain better performance than other baselines.
Overall, our proposed algorithm holds great promise for addressing this new variant of MCKP in the data-driven scenario. DDALS can serve as a baseline for future research. The code and benchmark sets will be open-sourced for further promoting research on DDCCMCKP. The algorithm framework can be further improved in two aspects. A new surrogate weight design may improve the efficiency and solution’s quality instead of the current design. From the other perspective, more effective filtering strategies that can balance the feasibility and optimality of the solutions could be explored in the future.

### v2

@@ C ET ECL RCL FSR Time C ET ECL RCL FSR Time
TABLE IX: Performance on the instances of APP. Due to the introduction of retransmission mechanisms, there is a significant difference in the results between LAB and APP.
Instance small $T_{\max}$ large $T_{\max}$
C ET ECL RCL FSR Time C ET ECL RCL FSR Time
ss2-Greedy 5.9$\pm$0.0 6.0$\pm$0.0 99.8$\pm$0.0 98.5$\pm$0.0 0 12.6$\pm$0.4 5.9$\pm$0.0 6.0$\pm$0.0 99.9$\pm$0.0 98.9$\pm$0.0 1 8.7$\pm$0.2
ss2-GA 5.3$\pm$0.1 303.2$\pm$2.3 99.2$\pm$0.0 97.3$\pm$0.5 0 4508.8$\pm$1198.4 5.3$\pm$0.1 272.9$\pm$2.2 99.5$\pm$0.1 98.1$\pm$0.2 0.03 3422.4$\pm$352.0
ss2-EDA 5.5$\pm$0.3 305.0$\pm$2.7 99.2$\pm$0.2 97.6$\pm$0.7 0.1 3323.2$\pm$1532.5 5.6$\pm$0.3 274.1$\pm$2.4 99.5$\pm$0.2 98.5$\pm$0.5 0.47 1185.3$\pm$1048.5
ss2-DDALS(O) 5.2$\pm$0.0 300.0$\pm$10.9 99.2$\pm$0.0 97.3$\pm$0.0 0 1010.5$\pm$98.7 5.2$\pm$0.0 297.5$\pm$7.8 99.6$\pm$0.0 98.1$\pm$0.0 0 862.2$\pm$78.7
ss2-DDALS(V1) 5.4$\pm$0.0 300.0$\pm$10.9 99.8$\pm$0.1 97.4$\pm$0.2 0 1010.5$\pm$98.7 5.2$\pm$0.0 297.5$\pm$7.8 99.6$\pm$0.0 98.1$\pm$0.0 0 862.2$\pm$78.7
ss2-DDALS(V2) 5.4$\pm$0.0 300.0$\pm$10.9 99.8$\pm$0.0 97.4$\pm$0.2 0 1010.5$\pm$98.7 5.4$\pm$0.1 297$\pm$7.8 99.9$\pm$0.1 98.2$\pm$0.2 0.1 862.2$\pm$78.7
ss2-DDALS(V3) 5.4$\pm$0.1 306 99.5$\pm$0.3 97.6$\pm$0.4 0 1086.0 5.5$\pm$0.1 298 99.7$\pm$0.2 98.5$\pm$0.3 0.5 858.0
ss4-Greedy 34.9$\pm$0.0 3.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1 0.2$\pm$0.0 28.5$\pm$0.0 4.0$\pm$0.0 100.0$\pm$0.0 99.0$\pm$0.0 1 0.3$\pm$0.0
ss4-GA 27.5$\pm$1.4 2008.9$\pm$7.9 99.8$\pm$0.2 90.3$\pm$0.5 0 251.8$\pm$24.5 20.5$\pm$2.7 1608.4$\pm$7.4 99.4$\pm$0.3 92.9$\pm$4.8 0 322.5$\pm$9.2
ss4-EDA 26.1$\pm$0.0 2005.7$\pm$2.3 100.0$\pm$0.0 90.1$\pm$0.0 0 333.9$\pm$11.8 18.3$\pm$2.7 1604.5$\pm$2.8 99.9$\pm$0.1 85.2$\pm$6.9 0 287.9$\pm$19.6
ss4-DDALS(O) 26.1$\pm$0.0 1988.9$\pm$27.8 100.0$\pm$0.0 90.1$\pm$0.0 0 33.3$\pm$2.3 16.6$\pm$0.8 1582.8$\pm$54.4 99.8$\pm$0.1 81.3$\pm$3.7 0 61.3$\pm$3.9
ss4-DDALS(V1) 26.1$\pm$0.0 1988.9$\pm$27.8 100.0$\pm$0.0 90.1$\pm$0.0 0 33.3$\pm$2.3 16.8$\pm$1.1 1582.8$\pm$54.4 99.8$\pm$0.1 81.8$\pm$4.3 0 61.3$\pm$3.9
ss4-DDALS(V2) 26.1$\pm$0.0 1988.9$\pm$27.8 100.0$\pm$0.0 90.1$\pm$0.0 0 33.3$\pm$2.3 18.3$\pm$2.2 1582.8$\pm$54.4 99.9$\pm$0.1 86.9$\pm$7.3 0 61.3$\pm$3.9
ss4-DDALS(V3) 29.7$\pm$2.5 1944 99.9$\pm$0.2 92.2$\pm$3.9 0.2 28.3 20.8$\pm$2.9 1586 99.7$\pm$0.3 88.3$\pm$6.3 0 59.3
ls2-Greedy 89.1$\pm$0.0 2.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 0.2$\pm$0.0 89.1$\pm$0.0 2.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1 0.3$\pm$0.
ls2-GA 76.7$\pm$2.8 22616.5$\pm$10.5 99.1$\pm$0.1 99.0$\pm$0.2 1 3925.1$\pm$155.5 60.9$\pm$2.7 21324.5$\pm$13.3 99.1$\pm$0.1 99.1$\pm$0.1 1 3382.2$\pm$64.9
ls2-EDA 79.3$\pm$1.1 22604.1$\pm$2.6 99.5$\pm$0.1 99.3$\pm$0.1 1 7174.3$\pm$44.2 59.9$\pm$0.8 21304.8$\pm$3.2 99.2$\pm$0.0 99.2$\pm$0.0 1 6970.8$\pm$88.9
ls2-DDALS(O) 72.0$\pm$0.3 22631.7$\pm$62.4 99.0$\pm$0.0 99.0$\pm$0.0 1 841.3$\pm$7.6 57.2$\pm$0.0 21265.2$\pm$73.3 99.3$\pm$0.1 99.3$\pm$0.0 1 776.4$\pm$6.6
ls2-DDALS(V1) 76.2$\pm$1.1 22631.7$\pm$62.4 99.5$\pm$0.1 99.5$\pm$0.1 1 841.3$\pm$7.6 59.0$\pm$1.2 21265.2$\pm$73.3 99.6$\pm$0.0 99.6$\pm$0.1 1 776.4$\pm$6.6
ls2-DDALS(V2) 73.2$\pm$0.4 22631.7$\pm$62.4 99.4$\pm$0.1 99.3$\pm$0.1 1 841.3$\pm$7.6 57.6$\pm$0.5 21265.2$\pm$73.3 99.4$\pm$0.1 99.5$\pm$0.1 1 776.4$\pm$6.6
ls2-DDALS(V3) 73.7$\pm$1.5 22536 99.2$\pm$0.1 99.0$\pm$0.2 1 846.2 58.3$\pm$0.6 21274 99.3$\pm$0.1 99.3$\pm$0.1 1 769.6
ls4-Greedy 238.1$\pm$0.0 3.0$\pm$0.0 99.4$\pm$0.0 99.0$\pm$0.0 1 1.5$\pm$0.0 193.7$\pm$0.0 15.0$\pm$0.0 99.2$\pm$0.0 98.8$\pm$0.0 1 7.0$\pm$0.2
ls4-GA 206.6$\pm$4.7 48018.6$\pm$13.4 99.0$\pm$0.0 98.6$\pm$0.3 0.73 46617.9$\pm$648.4 182.7$\pm$10.7 45015.2$\pm$11.7 99.0$\pm$0.0 98.5$\pm$0.2 0.6 42937.2$\pm$431.6
ls4-EDA 204.3$\pm$1.9 48004.5$\pm$3.5 99.0$\pm$0.0 99.0$\pm$0.0 1 45488.7$\pm$924.8 163.6$\pm$1.2 45003.5$\pm$2.4 99.1$\pm$0.1 98.8$\pm$0.1 1 42955.7$\pm$613.1
ls4-DDALS(O) 195.1$\pm$0.5 47999.9$\pm$83.0 99.0$\pm$0.0 98.9$\pm$0.0 1 16357.5$\pm$137.6 161.5$\pm$2.1 44440.5$\pm$344.6 99.0$\pm$0.0 98.5$\pm$0.3 0.4 15160.3$\pm$127.4
ls4-DDALS(V1) 209.2$\pm$2.8 47999.9$\pm$83.0 99.1$\pm$0.0 98.6$\pm$0.3 0.6 16357.5$\pm$137.6 169.4$\pm$4.2 44440.5$\pm$344.6 99.3$\pm$0.2 98.9$\pm$0.4 0.8 15160.3$\pm$127.4
ls4-DDALS(V2) 197.4$\pm$1.7 47999.9$\pm$83.0 99.1$\pm$0.0 99.0$\pm$0.1 1 16357.5$\pm$137.6 162.5$\pm$2.4 44440.5$\pm$344.6 99.1$\pm$0.1 98.6$\pm$0.2 0.7 15160.3$\pm$127.4
ls4-DDALS(V3) 196.8$\pm$0.8 48046 99.0$\pm$0.0 98.9$\pm$0.1 1 16280.8 163.5$\pm$3.5 44648 99.0$\pm$0.0 98.3$\pm$0.1 0.1 15162.6
Table VIII shows the results of ss3, ss4 and ls4 in the LAB. DDALS consistently obtains the lowest-cost solutions across all test instances. In contrast, Greedy requires fewer than 10 evaluations but yields the highest cost. The feasibility ratio for all algorithms is approximately 1, suggesting that our solution evaluation approach is able to achieve an evaluation close to the true confidence level. In comparison with other variations, the original version of DDALS consistently produces lower-cost solutions while maintaining a feasibility ratio of 1, demonstrating the accuracy of the solution evaluation approach.
@@
Table IX illustrates significant changes in the results of APP. In small-scale instance tests, only Greedy achieves a feasibility ratio of 1, while other algorithms nearly approach zero feasibility. Although DDALS obtain lower cost solutions than other algorithms, all of them are found to be infeasible. The primary challenge arises from the retransmission mechanism in APP, introducing discontinuities in item’s distribution that pose a considerable challenge to the solution evaluation approach. In cases of random sampling, small samples may overlook information about exceptional jumps, leading to a decline in solution evaluation performance. While Greedy achieves higher feasibility ratios, it sacrifices solution optimality. For large-scale problems, despite increased search difficulty with problem scale, DDALS consistently finds feasible solutions with the lowest cost. In the result of APP-ls4 with large $T_{\max}$, the feasibility ratio of DDALS(O) is 0.4, while that of DDALS(V1) is 0.8. Despite a 4.9% decrease in the average quality of solutions generated by DDALS(V1) relative to the original DDALS, the feasibility ratio increase by 100%. This emphasizes the effectiveness and necessity of filtering solutions in practical scenarios. Additionally, increasing the sample size $L$ is able to effectively enhances evaluation performance, capturing more information and improving the accuracy of solution evaluation approach. The results reveal that the runtime of DDALS is significantly shorter than both EDA and GA.
@@ VI-D Effectiveness of Components in DDALS
We conduct an ablation study to further assess the effectiveness of four components of DDALS, including CP, LSS, Degrade and FSS. Regarding the SFE, we can observe significant distinctions in the results between DDALS(O) and other variants in Table VIII and Table IX. To independently assess their effectiveness, four different variants of the original DDALS are tested on the instances of Lab:
@@
    r-CP DDALS: It uses random strategy for solution initialization instead of CP.
@@
    no-LSS DDALS: LSS is removed.
@@
    no-Degrade DDALS: Degrade is removed.
@@
    no-FSS DDALS: FSS is removed.
Table X presents the results. We compute the performance degradation ratio (PDR) by computing $(C_{2}-C_{1})/C1$ for each instance, where $C_{1}$ and $C_{2}$ are the average cost of solution obtained by the variants and original DDALS across 10 independent runs, respectively. The effect of the component is considered positive for a specific LAB instance when the PDR is greater than 0. The results indicate that none of the examined components adversely impacts the performance of DDALS. In most instances, these components enhance the performance of DDALS. Constructing utility for items proves beneficial for DDALS in achieving a favorable initial solution on average. While the purpose of LSS is to explore the solution space, a negative value on ls3 suggests that, due to large sample size, the algorithm may not converge within 30 iterations. Degrade contributes significantly to the performance of DDALS, addressing the problem by aiding its escape from local optima. FSS exhibits a slightly higher PDR than LSS. However, with the increase in the problem scale, the impact of FSS appears to diminish, indicating the need for more efficient and exploratory methods to address large-scale problems.
TABLE X: Average PDR on each instance of LAB. The highest PDR is highlighted through bolding and being underlined.
Instance r-CP no-LSS no-Degrade no-FSS
@@ VI-E Acceleration Performance of Accelerated Monte Carlo
The effectiveness of accelerated Monte Carlo (AMC) evaluation is confirmed through validation. We randomly selected 10,000 solutions of LAB-ls1 for assessment, employing both Monte Carlo (MC) and AMC methods. We record the runtime required for every 1000 evaluations, and the results are presented in Figure 6.
Experiment results reveal a nearly linear increase in the consumed time with the growth in the number of evaluation times. For various $T_{\max}$ settings, MC exhibits nearly constant time consumption, since the required number of simulations remains consistent across different $T_{\max}$. In contrast, AMC demonstrates superior acceleration, particularly with smaller $T_{\max}$. For example, compared to the MC’s runtime of $526.601s$, AMC completes the evaluation in $458.449s$ when $T_{\max}=11$, processing a fast-filtered solution count of 10,000. At $T_{\max}=23$, AMC takes approximately $471.395s$, with a pre-fast-filtered count of 9,718. As $T_{\max}$ increases, the challenge posed by the intensity of chance constraint diminishes, resulting in a higher percentage of feasible solutions in the entire solution space. When $T_{\max}=27$, AMC and MC take nearly the same time, while at $T_{\max}=35$, AMC needs $807.435s$, compared to approximately $500s$ for MC. The results show that accelerated mechanism may have a negative impact on the overall efficiency when the intensity of chance constraint is low. In conclusion, when implementing the accelerated mechanism, it is necessary to pre-assess whether the intensity of chance constraints is high or not.
Figure 6: Acceleration performance of AMC on LAB-ls1. The mechanism has a positive effect when the intensity of chance constraint is high, but has a negative impact on the overall efficiency as the intensity of chance constraint is lower, i.e., $T_{\max}$ becomes larger.
@@ VII Conclusion
In this work, motivated by practical applications, we study a novel variant of MCKP named chance-constrained MCKP. To address CCMCKP, we propose a data-driven adaptive local search algorithm. DDALS facilitates the data-driven solution evaluation approach that can effectively handle situations where the distributions of item weights are unknown, and incorporates a surrogate weight mechanism to ensure solution optimality. Moreover, a solution filtering mechanism is introduced to improve the feasibility of the output solutions. Two benchmark sets for CCMCKP have been established, one of which simulates the real business requirement of a 5G E2E network scenario for a communication company. Experimental results demonstrate the superiority of DDALS over other baselines on the two benchmarks. In cases with high intensity of chance constraints and limited amounts of sample data, DDALS still exhibits good performance. Ablation studies also confirm the effectiveness of each component of the algorithm.
Overall, our proposed algorithm holds great promise for addressing CCMCKP. Our proposed algorithm DDALS can serve as a baseline for future research. The source code of algorithm, benchmark sets and supplementary materials are available online for promoting further research on solving this problem. Furthermore, DDALS is not confined to CCMCKP but can be easily applied to other problems. Specifically, it is applicable to any chance-constrained problem that can be addressed via data-driven evaluation approach. While the algorithm framework is robust, it can be further improved in two aspects. A new surrogate weight design could enhance efficiency and solution quality beyond the current implementation. Additionally, it is necessary to further accelerate the evaluation process, since the evaluation of solutions is time-consuming due to extensive Monte Carlo simulations involved in each evaluation process.